OpenStudy (anonymous):

Sarah and Brian went shopping and spent a total of 47.50 brian spent 15.50 less than sarah how much did brian spend

OpenStudy (welshfella):

you could do it this way let brian spend x and Sarah spend y dollars so we have the equation x + y = 47.50

OpenStudy (welshfella):

and since Brian spent 15.50 less than Sarah y - x = 15.50

OpenStudy (welshfella):

now solve this system;- x + y = 47.50 y - x = 15.50

OpenStudy (welshfella):

Hint: add the 2 equations to get an equation in y only solve this to get what Sarah spent. Then find brians amount by subtracting 15.50 from Sarah's.

OpenStudy (anonymous):

answer is 16